# Bayesian nonparametric modeling of multivariate count data with an unknown number of traits

### v1

@@ 6 Analysis of the Infinito network
Consistent with the findings of Legramanti et al. (2022); Lu et al. (2025), the inferred group structures reveal intricate dynamics within this criminal organization. Notably, there is a clear tendency for clusters to form within individual locali, though the mechanisms governing group formation vary between affiliates and bosses. This pattern suggests a strategic design aimed at ensuring resilience across different levels of the organization. Furthermore, it highlights that while lower-ranked affiliates can form peripheral groups with relative ease, the formation of core groups, comprising all the bosses, is more constrained, as reflected in their progressively smaller sizes. A deeper examination of the cluster structure uncovers additional meaningful insights, further supporting observations made in Lu et al. (2025). For instance, juridical documents indicate that a specific affiliate from the blue locale was attempting to establish a new locale. This is reflected in our clustering results, where this individual is grouped primarily with lower-ranked members from the red locale (first cluster from the top), suggesting a deliberate strategy of recruiting new affiliates from that locale. Another striking pattern, even more evident in our analysis than in Lu et al. (2025), involves a cluster composed of bosses from multiple locali. Specifically, one green and one blue boss cluster together with two yellow bosses and two lower-ranked yellow affiliates. This aligns with reports that the green and blue bosses supported an unsuccessful attempt to increase the independence of the ‘Ndrangheta group in Lombardy from the ruling families in Calabria. Consequently, these individuals sought to distance themselves from their original locali and strengthen ties elsewhere. Our inferred clustering suggests that this shift primarily moved in the direction of the yellow locale, a hypothesis supported by the cluster composition. In contrast, Lu et al. (2025) infer this movement indirectly through node positioning in a distance-based network representation, rather than as a direct outcome of their methodology. A further distinction between our results and those in Lu et al. (2025) concerns the clustering of most yellow and purple lower-ranked affiliates. While our model merges them into a single cluster, their approach identifies them as distinct groups. However, their distance-based network representation highlights their close similarity. Finally, it is noteworthy that core groups of bosses frequently include a few lower-ranked affiliates, suggesting that these individuals may hold more central roles than indicated in the juridical documents. As observed in Lu et al. (2025), this is particularly evident in the case of a key affiliate grouped with the green locale bosses. In reality, this individual is a high-ranking member with crucial coordinating responsibilities across multiple locali.
@@
A key property of both the known-groups and unknown-groups models is that they explicitly allow for the estimation of the number of meetings that went undetected by investigators. Notably, such an analysis cannot be conducted in Legramanti et al. (2022); Lu et al. (2025), as those models implicitly assume a fixed number of meetings. In the analysis of the Infinito criminal data, we rely on the inference provided by the unknown-groups model, which was identified as the better-fitting specification based on the comparisons discussed above. Our unknown-groups model suggests the likely presence of a few number of undetected meetings, which aligns with expectations and confirms the need of accounting for unseen binary traits. Figure 4(a) displays the posterior distribution of the number of these unseen meetings, providing a probabilistic estimate of their occurrence. This information may provide guidance to law enforcements to understand the amount of meetings that went undetected. Moreover, as formally proved in Section 4.2, accounting for unseen meetings directly affects the clustering structure and, indirectly, the estimated adjacency matrices in Figure 3. The overall good fit of the proposed unknown-groups model to data, highlighted in Figure 3, also validates the estimate of the number of unseen meetings, which results to be much smaller than the estimate provided under the known-groups setting using the clustering induced by locali membership, reported in Figure 4(b). We also compare these findings with those obtained from the negative binomial mixture of bbs (Ghilotti et al., 2025), the natural benchmark for estimating the number of unseen meetings. Hyperparameters are set as in Section S4.2. By construction, this model imposes a homogeneous exchangeable structure, i.e. it does not take into account any group structure, leading to posterior expectations of the adjacency matrix that assign identical values across all entries. As a result, it fails to reproduce the heterogeneity observed in the Infinito criminal data (Figure 3(a)). Conseqeuntly, its inference on the number of unseen meetings $N^{\prime}$ substantially differs from that of our model. In particular, Figure 4(c) shows that the competitor places high posterior probability on the absence of undetected meetings, in contrast with the inference obtained under the unknown-groups model (Figure 4(a)). Model comparison via the waic further supports the data-driven clustering approach. Indeed, the negative binomial mixture of bbs attains a waic of 5392-5392, which is worse not only than the unknown-groups model but also than the known-groups specification.
@@ 7 Discussion
Although we focused on the analysis of criminal networks, the scope of our approach extends beyond this domain. For instance, we plan to explore its use to cluster microbiome profiles of different subjects, where traits correspond to operational taxonomic units observed with different frequencies in each subject. The importance of trait allocation models for microbiome data analysis has been recently emphasized by James et al. (2025). Another promising area of application is ecology, where our method can be used to cluster sites that exhibit similar patterns in terms of species occurrences (Ghilotti et al., 2025). In this context the traits are the species, observed with varying frequencies in each site.
In addition, our paper also serves as a springboard for future methodological developments. First, all results in Section S2 of the Supplementary Material hold true for general crvs, offering a theoretical basis for a broad class of models and providing practitioners with new modeling opportunities. Second, the results in Section 3 are the basic building blocks for addressing extrapolation problems involving dependent populations of traits. For example, one can estimate the number of unseen traits in an additional sample of subjects, thereby extending the results of Masoero et al. (2022); Camerlenghi et al. (2024); Ghilotti et al. (2025) when multiple-sample data are available. While Shen et al. (2024) focused on predicting new genetic variants in presence of two dependent populations, our approach offers additional flexibility in ecological applications by allowing for finite species richness, since the number of traits is assumed finite but random. Finally, we believe that it is not difficult to extend our framework to cluster a collection of trait allocations, rather than individual subjects. This would lead to models akin to nested processes (Rodríguez et al., 2008) within the trait allocation context. The aforementioned applications and methodological developments are left for future research.
@@ Acknowledgments
In this section, we provide an account on completely random vectors (crvs), with the goal of illustrating the connection between the processes presented in Section 3.1 and completely random vectors. The class of crvs can be considered as a generalization of the class of completely random measures (crms), introduced by Kingman (1967), to the multivariate setting and played a leading role in many successful stories in the bnp literature. Indeed many authors have focused on crvs to model the prior opinion in presence of partially exchangeable data, i.e., data organized in groups (Griffin and Leisen, 2017; Lijoi et al., 2014; Camerlenghi et al., 2019). The use of crvs also allows the quantification of dependence induced by the prior across the different groups of observations (Catalano et al., 2021, 2024).
To mathematically introduce the class of crvs, consider a Polish space $\mathds{X}$ equipped with the corresponding Borel $\sigma$-algebra $\mathscr{X}$; consistently with the rest of the paper, $\mathds{X}$ represents the space of trait labels. Denote by $(\mathsf{M}_{\mathds{X}},\mathcal{M}_{\mathds{X}})$ the space of boundedly finite (non-negative) measures on $\mathds{X}$ equipped with the corresponding Borel $\sigma$-algebra generated by the topology of weak#-convergence (Daley and Vere-Jones, 2008). We also indicate by $\mathsf{M}_{\mathds{X}}^{d}$ the $d$-fold product space of $\mathsf{M}_{\mathds{X}}$, which is naturally endowed with the product $\sigma$-algebra. Remind that a random vector of measures $\tilde{\bm{\mu}}=(\tilde{\mu}_{1},\ldots,\tilde{\mu}_{d})$ is a measurable function from a suitable probability space, say $(\Omega,\mathscr{A},\mathds{P})$, taking values in $(\mathsf{M}_{\mathds{X}}^{d},\mathcal{M}_{\mathds{X}}^{d})$. In the sequel, we will use the notation $\tilde{\bm{\mu}}(A)=(\tilde{\mu}_{1}(A),\ldots,\tilde{\mu}_{d}(A))$ to indicate the random vector evaluated on a set $A\in\mathscr{X}$. Then, crvs are defined as follows.
@@
See Kallenberg (2017, Theorem 3.19). As a common practice in bnp models, here we remove the deterministic drift and the part with fixed locations, thus working only with component (iii); in other words we consider crvs of the type:
@@
Similarly to the crm case, any crv can be expressed as a functional of a Poisson point process (Baccelli et al., 2020) on the space $(0,\infty)^{d}\times\mathds{X}$. Indeed, consider the Laplace functional of $\tilde{\bm{\mu}}$, which is defined as
@@
for any bounded Borel set $B\in\mathscr{X}$, where $\bm{\theta}||$ denotes the Euclidean norm of the vector $\bm{\theta}=(\theta_{1},\ldots,\theta_{d})$. In the present paper we consider homogeneous crvs, namely crvs whose Lévy intensity satisfies $\nu(\mathrm{d}\theta_{1}\,\ldots\,\mathrm{d}\theta_{d}\,\mathrm{d}x)=\rho_{d}(\mathrm{d}\theta_{1}\,\ldots\,\mathrm{d}\theta_{d})\alpha(\mathrm{d}x)$. In particular, assume that $\alpha(\,\cdot\,)$ is a finite non-atomic measure on $\mathds{X}$ with total mass $\lambda$, thus set $P_{0}(\,\cdot\,):=\alpha(\,\cdot\,)/\lambda$. In the sequel, we use the notation $\tilde{\bm{\mu}}\sim\textsc{crv}(\rho_{d},\lambda,P_{0})$ to indicate the distribution of a homogeneous crv. We refer to Kallenberg (2017) for additional details on crvs.
@@
When $\tilde{\bm{\mu}}$ is a fcrv, we write $\tilde{\bm{\mu}}\sim\textsc{fcrv}(\rho_{d},\lambda,P_{0})$, where $\rho_{d}$ is necessarily finite. Without loss of generality, assume that $\rho_{d}=H^{(d)}$ is a probability measure and adjust $\lambda$ to account for it. As we will discuss in the next proposition, the specification finite crv is motivated by the fact that fcrvs can be seen as functionals of finite Poisson point processes (Baccelli et al., 2020, Section 4.3). In particular, the random measures $\tilde{\mu}_{1},\ldots,\tilde{\mu}_{d}$ are supported on a (finite) Poisson random number of common atoms.
@@
In addition to the example of fcrvs discussed in the main paper, there are many other tractable classes of crvs that can be analyzed based on the general theory of the present section, and that are of potential interest in applications. For example, Theorems S1-S2-S3 may be applied to obtain marginal, posterior, and predictive distributions of additive crms introduced by Lijoi et al. (2014). Another example of interest that can be addressed with the provided theory relates to compound random measures (Griffin and Leisen, 2017).
@@
Before moving to more structured simulation studies, we first provide empirical evidence on the discrepancy highlighted in Proposition 1, which may indeed lead to substantial inferential differences. To this end, we consider a simulated binary-outcomes example, where we compare the inference obtained from our proposed model (16) with that of a latent class model in which $N=k$ is fixed, thus disregarding the possibility of unseen traits, as in Dunson and Xing (2009). We consider data organized into $d=5$ groups, with a total of $N=500$ traits that may potentially be observed. Group-specific probability vectors $\theta_{jq}$ are generated by drawing iid values from a $\text{Beta}(0.1,10)$ distribution, for $q=1,\ldots,5$ and $j=1,\ldots,500$. Subject-specific binary vectors are then obtained by independently sampling Bernoulli random variables $\tilde{A}_{ijq}$ with success probabilities $\theta_{jq}$, determined by the group allocation. Samples are drawn from the five groups with different sizes, namely $n_{1}=100$, $n_{2}=60$, $n_{3}=40$, $n_{4}=20$, and $n_{5}=20$, where $n_{q}$ denotes the sample size of group $q$. In the resulting dataset, $K_{n}=k=293$ traits are observed out of the total $N=500$.
@@
Figure S1: Posterior distribution of the number of clusters under model (16) with binary traits (in skyblue) and Dunson and Xing (2009) model (in orange). The true number of clusters in the generating mechanism is $5$.
@@
This section evaluates the performance of the latent class model with unseen traits described in Section 4, referred to as unknown-groups model. Motivated by our focus on criminal network analysis, we generate synthetic datasets designed to capture a realistically complex structure. The analysis of criminal networks (Legramanti et al., 2022; Lu et al., 2025) typically aims at detecting groups of affiliates who share similar connectivity patterns. These patterns are encoded in a weighted adjacency matrix, which records the frequency of interactions between pairs of individuals. An interaction is defined as the co-attendance of two affiliates at the same meeting. The weighted adjacency matrix thus summarizes the raw data, originally collected as multivariate binary observations, where each affiliate is associated with the list of meetings they attended. Formally, if the dataset comprises $n$ affiliates and $k$ distinct meetings, the attendance information may be arranged in an $n\times k$ binary matrix $\bm{A}$, with entries $A_{i\ell}$ indicating whether individual $i$ attended meeting $\ell$. The corresponding observed weighted adjacency matrix, denoted by $\bm{W}$, is then obtained as $\bm{W}=\bm{A}\bm{A}^{T}$. Its generic entry $W_{ii^{\prime}}$ represents the number of meetings jointly attended by affiliates $i$ and $i^{\prime}$, that is, $W_{ii^{\prime}}=\sum_{\ell=1}^{k}A_{i\ell}A_{i^{\prime}\ell}$ for $i\neq i^{\prime}$. Our framework induces a probabilistic structure on $\bm{W}$: each entry $W_{ii^{\prime}}$ can be interpreted as the realization of a sum of Bernoulli random variables. To assess the flexibility of the proposed methodology in the analysis of criminal networks, we consider two simulated scenarios, each characterized by distinct properties of the resulting weighted adjacency matrix. These scenarios are constructed to produce structures akin to those investigated in Lu et al. (2025), thereby closely reflecting patterns observed in real-world criminal networks. A detailed description of the two scenarios is provided below.
@@
We benchmark our approach against the negative binomial mixture of bbs introduced in Ghilotti et al. (2025), which represents the natural competitor for estimating the number of unseen traits. This model, however, is specifically tailored to homogeneous exchangeable settings. It falls within the class described in Example 1, corresponding to the binary traits specification of model (4) with prior given in equations (5)–(6), under the special case $d=1$. In this framework, $H(\cdot;\psi)$ is the beta distribution with parameters $(-\alpha,\alpha+\beta)$. The beta parameters $(a,b)=(-\alpha,\alpha+\beta)$ are equipped with independent gamma priors with hyperparameters $(\alpha_{a},\beta_{a})$ and $(\alpha_{b},\beta_{b})$, respectively. The model parameter $\lambda$, governing the Poisson-distributed total number of meetings, is assigned a gamma prior with parameters $(\alpha_{\lambda},\beta_{\lambda})$, as detailed in Section 3.3. By construction, the negative binomial mixture of bbs enforces homogeneity, resulting in posterior expectations for $\bm{W}$ that are constant across entries. Consequently, the model cannot capture potential heterogeneity in connectivity patterns that may emerge from the observed adjacency matrix. For all comparisons presented in this paper, hyperparameters are fixed as in Ghilotti et al. (2025), with the prior variance for $N$ set to be comparable to that used in the proposed unknown-groups model for each application. Posterior inference relies on 100,000 iterations, with the first 10,000 samples discarded as burn-in and a thinning interval of 5.
@@
We further compare our methodology to the negative binomial mixture of bbs introduced in Ghilotti et al. (2025). Both models allow inference on the number of unseen meetings $N^{\prime}$, which in the framework of Theorem 2 corresponds to the number of atoms $N^{\prime}$ in the measures $\mu_{q}^{\prime}$. Figure S3 reports the posterior distributions of $N^{\prime}$ for the two models. Our proposed model provides an accurate estimate, with uncertainty appropriately quantified. In contrast, the negative binomial mixture of bbs substantially overestimates $N^{\prime}$, failing to address the unseen features problem in this heterogeneous setting.
@@
Figure S3: Synthetic network data, Scenario 1. Posterior distribution of the number of unseen meetings under (a) the unknown-groups model and (b) the negative binomial mixture of bbs in Ghilotti et al. (2025). The vertical red dotted lines denote the true number of unseen meetings.
@@
We next examine inference on the number of unseen meetings $N^{\prime}$. Figure S4 presents the posterior distribution of $N^{\prime}$ under the proposed unknown-groups model and the negative binomial mixture of bbs in Ghilotti et al. (2025). Figure 4(a) shows that our unknown-groups model recovers the number of unseen meetings accurately, while also providing a coherent quantification of uncertainty. At first glance, the competitor appears to perform well, as its posterior distribution is centered closer to the true value (red vertical line), albeit with substantially greater dispersion. However, this apparent accuracy is largely coincidental. In fact, since the negative binomial mixture of bbs enforces homogeneity, it cannot accommodate the heterogeneity clearly visible in the simulated adjacency matrix (Figure 2(b)). For ease of comparison, Figure S5 juxtaposes the simulated adjacency matrix, the posterior expectation of the adjacency matrix under our unknown-groups model, and the posterior expectation of the adjacency matrix under the competitor. The figure highlights that while our model reproduces the complex connectivity structure, the competitor collapses to an unrealistic homogeneous pattern.
@@
Figure S4: Synthetic network data, Scenario 2. Posterior distribution of the number of unseen meetings under (a) the unknown-groups model and (b) the negative binomial mixture of bbs in Ghilotti et al. (2025). The vertical red dotted lines denote the true number of unseen meetings.

### v2

@@ 6 Analysis of the Infinito network
Consistent with the findings of esbm_rigon; Lu2025, the inferred group structures reveal intricate dynamics within this criminal organization. Notably, there is a clear tendency for clusters to form within individual locali, though the mechanisms governing group formation vary between affiliates and bosses. This pattern suggests a strategic design aimed at ensuring resilience across different levels of the organization. Furthermore, it highlights that while lower-ranked affiliates can form peripheral groups with relative ease, the formation of core groups, comprising all the bosses, is more constrained, as reflected in their progressively smaller sizes. A deeper examination of the cluster structure uncovers additional meaningful insights, further supporting observations made in Lu2025. For instance, juridical documents indicate that a specific affiliate from the blue locale was attempting to establish a new locale. This is reflected in our clustering results, where this individual is grouped primarily with lower-ranked members from the red locale (first cluster from the top), suggesting a deliberate strategy of recruiting new affiliates from that locale. Another striking pattern, even more evident in our analysis than in Lu2025, involves a cluster composed of bosses from multiple locali. Specifically, one green and one blue boss cluster together with two yellow bosses and two lower-ranked yellow affiliates. This aligns with reports that the green and blue bosses supported an unsuccessful attempt to increase the independence of the ‘Ndrangheta group in Lombardy from the ruling families in Calabria. Consequently, these individuals sought to distance themselves from their original locali and strengthen ties elsewhere. Our inferred clustering suggests that this shift primarily moved in the direction of the yellow locale, a hypothesis supported by the cluster composition. In contrast, Lu2025 infer this movement indirectly through node positioning in a distance-based network representation, rather than as a direct outcome of their methodology. A further distinction between our results and those in Lu2025 concerns the clustering of most yellow and purple lower-ranked affiliates. While our model merges them into a single cluster, their approach identifies them as distinct groups. However, their distance-based network representation highlights their close similarity. Finally, it is noteworthy that core groups of bosses frequently include a few lower-ranked affiliates, suggesting that these individuals may hold more central roles than indicated in the juridical documents. As observed in Lu2025, this is particularly evident in the case of a key affiliate grouped with the green locale bosses. In reality, this individual is a high-ranking member with crucial coordinating responsibilities across multiple locali.
@@
A key property of both the known-groups and unknown-groups models is that they explicitly allow for the estimation of the number of meetings that went undetected by investigators. Notably, such an analysis cannot be conducted in esbm_rigon; Lu2025, as those models implicitly assume a fixed number of meetings. In the analysis of the Infinito criminal data, we rely on the inference provided by the unknown-groups model, which was identified as the better-fitting specification based on the comparisons discussed above. Our unknown-groups model suggests the likely presence of a few number of undetected meetings, which aligns with expectations and confirms the need of accounting for unseen binary traits. Figure 4(a) displays the posterior distribution of the number of these unseen meetings, providing a probabilistic estimate of their occurrence. This information may provide guidance to law enforcements to understand the amount of meetings that went undetected. Moreover, as formally proved in Section 4.2, accounting for unseen meetings directly affects the clustering structure and, indirectly, the estimated adjacency matrices in Figure 3. The overall good fit of the proposed unknown-groups model to data, highlighted in Figure 3, also validates the estimate of the number of unseen meetings, which results to be much smaller than the estimate provided under the known-groups setting using the clustering induced by locali membership, reported in Figure 4(b). We also compare these findings with those obtained from the negative binomial mixture of bbs (Ghilotti2025), the natural benchmark for estimating the number of unseen meetings. Hyperparameters are set as in Section S4.2. By construction, this model imposes a homogeneous exchangeable structure, i.e. it does not take into account any group structure, leading to posterior expectations of the adjacency matrix that assign identical values across all entries. As a result, it fails to reproduce the heterogeneity observed in the Infinito criminal data (Figure 3(a)). Conseqeuntly, its inference on the number of unseen meetings $N^{\prime}$ substantially differs from that of our model. In particular, Figure 4(c) shows that the competitor places high posterior probability on the absence of undetected meetings, in contrast with the inference obtained under the unknown-groups model (Figure 4(a)). Model comparison via the waic further supports the data-driven clustering approach. Indeed, the negative binomial mixture of bbs attains a waic of 5392-5392, which is worse not only than the unknown-groups model but also than the known-groups specification.
@@ 7 Discussion
Although we focused on the analysis of criminal networks, the scope of our approach extends beyond this domain. For instance, we plan to explore its use to cluster microbiome profiles of different subjects, where traits correspond to operational taxonomic units observed with different frequencies in each subject. The importance of trait allocation models for microbiome data analysis has been recently emphasized by james2025microbiome. Another promising area of application is ecology, where our method can be used to cluster sites that exhibit similar patterns in terms of species occurrences (Ghilotti2025). In this context the traits are the species, observed with varying frequencies in each site.
In addition, our paper also serves as a springboard for future methodological developments. First, all results in Section S2 of the Supplementary Material hold true for general crvs, offering a theoretical basis for a broad class of models and providing practitioners with new modeling opportunities. Second, the results in Section 3 are the basic building blocks for addressing extrapolation problems involving dependent populations of traits. For example, one can estimate the number of unseen traits in an additional sample of subjects, thereby extending the results of Mas22; Camerlenghi2024; Ghilotti2025 when multiple-sample data are available. While shen2024 focused on predicting new genetic variants in presence of two dependent populations, our approach offers additional flexibility in ecological applications by allowing for finite species richness, since the number of traits is assumed finite but random. Finally, we believe that it is not difficult to extend our framework to cluster a collection of trait allocations, rather than individual subjects. This would lead to models akin to nested processes (Rod(08)) within the trait allocation context. The aforementioned applications and methodological developments are left for future research.
@@ Acknowledgments
In this section, we provide an account on completely random vectors (crvs), with the goal of illustrating the connection between the processes presented in Section 3.1 and completely random vectors. The class of crvs can be considered as a generalization of the class of completely random measures (crms), introduced by Kingman(67), to the multivariate setting and played a leading role in many successful stories in the bnp literature. Indeed many authors have focused on crvs to model the prior opinion in presence of partially exchangeable data, i.e., data organized in groups (Gri(17); Nip(14); Cam(19)AoS). The use of crvs also allows the quantification of dependence induced by the prior across the different groups of observations (Cat(21)AoS; Lav(21)).
To mathematically introduce the class of crvs, consider a Polish space $\mathds{X}$ equipped with the corresponding Borel $\sigma$-algebra $\mathscr{X}$; consistently with the rest of the paper, $\mathds{X}$ represents the space of trait labels. Denote by $(\mathsf{M}_{\mathds{X}},\mathcal{M}_{\mathds{X}})$ the space of boundedly finite (non-negative) measures on $\mathds{X}$ equipped with the corresponding Borel $\sigma$-algebra generated by the topology of weak#-convergence (Daley_volII). We also indicate by $\mathsf{M}_{\mathds{X}}^{d}$ the $d$-fold product space of $\mathsf{M}_{\mathds{X}}$, which is naturally endowed with the product $\sigma$-algebra. Remind that a random vector of measures $\tilde{\bm{\mu}}=(\tilde{\mu}_{1},\ldots,\tilde{\mu}_{d})$ is a measurable function from a suitable probability space, say $(\Omega,\mathscr{A},\mathds{P})$, taking values in $(\mathsf{M}_{\mathds{X}}^{d},\mathcal{M}_{\mathds{X}}^{d})$. In the sequel, we will use the notation $\tilde{\bm{\mu}}(A)=(\tilde{\mu}_{1}(A),\ldots,\tilde{\mu}_{d}(A))$ to indicate the random vector evaluated on a set $A\in\mathscr{X}$. Then, crvs are defined as follows.
@@
See Kallenberg_2017. As a common practice in bnp models, here we remove the deterministic drift and the part with fixed locations, thus working only with component (iii); in other words we consider crvs of the type:
@@
Similarly to the crm case, any crv can be expressed as a functional of a Poisson point process (baccelli) on the space $(0,\infty)^{d}\times\mathds{X}$. Indeed, consider the Laplace functional of $\tilde{\bm{\mu}}$, which is defined as
@@
for any bounded Borel set $B\in\mathscr{X}$, where $\bm{\theta}||$ denotes the Euclidean norm of the vector $\bm{\theta}=(\theta_{1},\ldots,\theta_{d})$. In the present paper we consider homogeneous crvs, namely crvs whose Lévy intensity satisfies $\nu(\mathrm{d}\theta_{1}\,\ldots\,\mathrm{d}\theta_{d}\,\mathrm{d}x)=\rho_{d}(\mathrm{d}\theta_{1}\,\ldots\,\mathrm{d}\theta_{d})\alpha(\mathrm{d}x)$. In particular, assume that $\alpha(\,\cdot\,)$ is a finite non-atomic measure on $\mathds{X}$ with total mass $\lambda$, thus set $P_{0}(\,\cdot\,):=\alpha(\,\cdot\,)/\lambda$. In the sequel, we use the notation $\tilde{\bm{\mu}}\sim\textsc{crv}(\rho_{d},\lambda,P_{0})$ to indicate the distribution of a homogeneous crv. We refer to Kallenberg_2017 for additional details on crvs.
@@
When $\tilde{\bm{\mu}}$ is a fcrv, we write $\tilde{\bm{\mu}}\sim\textsc{fcrv}(\rho_{d},\lambda,P_{0})$, where $\rho_{d}$ is necessarily finite. Without loss of generality, assume that $\rho_{d}=H^{(d)}$ is a probability measure and adjust $\lambda$ to account for it. As we will discuss in the next proposition, the specification finite crv is motivated by the fact that fcrvs can be seen as functionals of finite Poisson point processes (baccelli, Section 4.3). In particular, the random measures $\tilde{\mu}_{1},\ldots,\tilde{\mu}_{d}$ are supported on a (finite) Poisson random number of common atoms.
@@
In addition to the example of fcrvs discussed in the main paper, there are many other tractable classes of crvs that can be analyzed based on the general theory of the present section, and that are of potential interest in applications. For example, Theorems S1-S2-S3 may be applied to obtain marginal, posterior, and predictive distributions of additive crms introduced by Nip(14). Another example of interest that can be addressed with the provided theory relates to compound random measures (Gri(17)).
@@
Before moving to more structured simulation studies, we first provide empirical evidence on the discrepancy highlighted in Proposition 1, which may indeed lead to substantial inferential differences. To this end, we consider a simulated binary-outcomes example, where we compare the inference obtained from our proposed model (16) with that of a latent class model in which $N=k$ is fixed, thus disregarding the possibility of unseen traits, as in Dunson2009. We consider data organized into $d=5$ groups, with a total of $N=500$ traits that may potentially be observed. Group-specific probability vectors $\theta_{jq}$ are generated by drawing iid values from a $\text{Beta}(0.1,10)$ distribution, for $q=1,\ldots,5$ and $j=1,\ldots,500$. Subject-specific binary vectors are then obtained by independently sampling Bernoulli random variables $\tilde{A}_{ijq}$ with success probabilities $\theta_{jq}$, determined by the group allocation. Samples are drawn from the five groups with different sizes, namely $n_{1}=100$, $n_{2}=60$, $n_{3}=40$, $n_{4}=20$, and $n_{5}=20$, where $n_{q}$ denotes the sample size of group $q$. In the resulting dataset, $K_{n}=k=293$ traits are observed out of the total $N=500$.
@@
Figure S1: Posterior distribution of the number of clusters under model (16) with binary traits (in skyblue) and Dunson2009 model (in orange). The true number of clusters in the generating mechanism is $5$.
@@
This section evaluates the performance of the latent class model with unseen traits described in Section 4, referred to as unknown-groups model. Motivated by our focus on criminal network analysis, we generate synthetic datasets designed to capture a realistically complex structure. The analysis of criminal networks (esbm_rigon; Lu2025) typically aims at detecting groups of affiliates who share similar connectivity patterns. These patterns are encoded in a weighted adjacency matrix, which records the frequency of interactions between pairs of individuals. An interaction is defined as the co-attendance of two affiliates at the same meeting. The weighted adjacency matrix thus summarizes the raw data, originally collected as multivariate binary observations, where each affiliate is associated with the list of meetings they attended. Formally, if the dataset comprises $n$ affiliates and $k$ distinct meetings, the attendance information may be arranged in an $n\times k$ binary matrix $\bm{A}$, with entries $A_{i\ell}$ indicating whether individual $i$ attended meeting $\ell$. The corresponding observed weighted adjacency matrix, denoted by $\bm{W}$, is then obtained as $\bm{W}=\bm{A}\bm{A}^{T}$. Its generic entry $W_{ii^{\prime}}$ represents the number of meetings jointly attended by affiliates $i$ and $i^{\prime}$, that is, $W_{ii^{\prime}}=\sum_{\ell=1}^{k}A_{i\ell}A_{i^{\prime}\ell}$ for $i\neq i^{\prime}$. Our framework induces a probabilistic structure on $\bm{W}$: each entry $W_{ii^{\prime}}$ can be interpreted as the realization of a sum of Bernoulli random variables. To assess the flexibility of the proposed methodology in the analysis of criminal networks, we consider two simulated scenarios, each characterized by distinct properties of the resulting weighted adjacency matrix. These scenarios are constructed to produce structures akin to those investigated in Lu2025, thereby closely reflecting patterns observed in real-world criminal networks. A detailed description of the two scenarios is provided below.
@@
We benchmark our approach against the negative binomial mixture of bbs introduced in Ghilotti2025, which represents the natural competitor for estimating the number of unseen traits. This model, however, is specifically tailored to homogeneous exchangeable settings. It falls within the class described in Example 1, corresponding to the binary traits specification of model (4) with prior given in equations (5)–(6), under the special case $d=1$. In this framework, $H(\cdot;\psi)$ is the beta distribution with parameters $(-\alpha,\alpha+\beta)$. The beta parameters $(a,b)=(-\alpha,\alpha+\beta)$ are equipped with independent gamma priors with hyperparameters $(\alpha_{a},\beta_{a})$ and $(\alpha_{b},\beta_{b})$, respectively. The model parameter $\lambda$, governing the Poisson-distributed total number of meetings, is assigned a gamma prior with parameters $(\alpha_{\lambda},\beta_{\lambda})$, as detailed in Section 3.3. By construction, the negative binomial mixture of bbs enforces homogeneity, resulting in posterior expectations for $\bm{W}$ that are constant across entries. Consequently, the model cannot capture potential heterogeneity in connectivity patterns that may emerge from the observed adjacency matrix. For all comparisons presented in this paper, hyperparameters are fixed as in Ghilotti2025, with the prior variance for $N$ set to be comparable to that used in the proposed unknown-groups model for each application. Posterior inference relies on 100,000 iterations, with the first 10,000 samples discarded as burn-in and a thinning interval of 5.
@@
We further compare our methodology to the negative binomial mixture of bbs introduced in Ghilotti2025. Both models allow inference on the number of unseen meetings $N^{\prime}$, which in the framework of Theorem 2 corresponds to the number of atoms $N^{\prime}$ in the measures $\mu_{q}^{\prime}$. Figure S3 reports the posterior distributions of $N^{\prime}$ for the two models. Our proposed model provides an accurate estimate, with uncertainty appropriately quantified. In contrast, the negative binomial mixture of bbs substantially overestimates $N^{\prime}$, failing to address the unseen features problem in this heterogeneous setting.
@@
Figure S3: Synthetic network data, Scenario 1. Posterior distribution of the number of unseen meetings under (a) the unknown-groups model and (b) the negative binomial mixture of bbs in Ghilotti2025. The vertical red dotted lines denote the true number of unseen meetings.
@@
We next examine inference on the number of unseen meetings $N^{\prime}$. Figure S4 presents the posterior distribution of $N^{\prime}$ under the proposed unknown-groups model and the negative binomial mixture of bbs in Ghilotti2025. Figure 4(a) shows that our unknown-groups model recovers the number of unseen meetings accurately, while also providing a coherent quantification of uncertainty. At first glance, the competitor appears to perform well, as its posterior distribution is centered closer to the true value (red vertical line), albeit with substantially greater dispersion. However, this apparent accuracy is largely coincidental. In fact, since the negative binomial mixture of bbs enforces homogeneity, it cannot accommodate the heterogeneity clearly visible in the simulated adjacency matrix (Figure 2(b)). For ease of comparison, Figure S5 juxtaposes the simulated adjacency matrix, the posterior expectation of the adjacency matrix under our unknown-groups model, and the posterior expectation of the adjacency matrix under the competitor. The figure highlights that while our model reproduces the complex connectivity structure, the competitor collapses to an unrealistic homogeneous pattern.
@@
Figure S4: Synthetic network data, Scenario 2. Posterior distribution of the number of unseen meetings under (a) the unknown-groups model and (b) the negative binomial mixture of bbs in Ghilotti2025. The vertical red dotted lines denote the true number of unseen meetings.